# Differential Privacy for Deep Learning in Medicine

### v1

@@ 5.1 Overview of Application Domains
Kaess et al. [71] CT Centralized Detecting CT contrast agents GM/NA
Khanna et al. [72] Genomic data FL Classification GM/RDP
Lal et al. [73] Tabular data Centralized Classification of fetal status GM/MA
Mehmood et al. [74] MRI FL Classification (brain tumor, AD) GM/NA
Mueller et al. [75] Image-derived graph Centralized Whole-graph classification GM/RDP
Odeyomi et al. [76] Chest X-ray Centralized Pneumonia classification GM/RDP
P3SGD [77] Histopathology Centralized Classification GM/MA
Pan et al. [78] Histopathology FL Segmentation GM/NA
PP-LDG [79] CT Centralized Segmentation (liver tumors) GM/MA
Shiri et al. [80] Chest CT FL COVID-19 prognosis GM/NA
Tayebi Arasteh et al. [13] X-ray Centralized Classification GM/RDP
Tayebi Arasteh et al. [5] X-ray, abdominal CT Centralized Classification GM/RDP
Tayebi Arasteh et al. [81] Pathological speech Centralized Neurological disorder analysis GM/RDP
Torfi et al. [82] EEG, EHR, ECG, biopsy Centralized Data generation GM/RDP
Wang et al. [83] Tabular data FL Classification GM/MA
Ziller et al. [84] X-ray, CT, fundoscopy Centralized Classification, segmentation GM/GDP
@@
5.1.1 Modalities and Clinical Contexts
The reviewed studies apply DP-SGD across a broad range of medical data types, with imaging emerging as the most common modality. Chest X-rays dominate classification tasks [13, 5, 64, 67, 84, 65, 79, 71], while MRI—particularly brain and cardiac cine MRI—is used in both classification and generative modeling applications [74, 62]. Histopathology appears in several studies focusing on cancer classification and segmentation [58, 78, 77, 60].
Beyond imaging, DP-SGD has been applied to structured data such as EHRs and tabular datasets [66, 73, 61, 59], genomic sequences [63, 72], biosignals including EEG and ECG [82], and pathological speech [81]. Time series data is also explored in classification contexts [70], and some generative frameworks integrate multimodal data such as EEG, biopsy, and EHR [82].
@@ 5.1.1 Modalities and Clinical Contexts
Classification remains the predominant downstream task [13, 5, 64, 67, 84, 65, 71, 74, 60, 66, 73, 72, 68, 70], followed by generative modeling [62, 61, 59, 82]. Despite growing interest, applications involving longitudinal data or multimodal fusion remain underexplored.
5.1.2 Deployment Settings: Centralized vs Federated Learning
DP-SGD is employed in both centralized [13, 5, 64, 63, 66, 84, 75, 76, 73, 68, 79, 77, 62, 81, 61, 59, 60, 71, 82, 70] and federated [80, 83, 67, 72, 69, 58, 74, 65, 78] training pipelines. In centralized settings, models are trained on institutional data and privatized using DP-SGD, protecting against post-deployment leakage. These include imaging [13, 5, 64, 84], structured/tabular data [66, 73], and EHRs [61, 59].
@@ 5.1.2 Deployment Settings: Centralized vs Federated Learning
FL-based applications involve multi-institutional collaboration, where local clients update models on-site and share only privatized gradients. These setups target imaging [80, 67, 69], genomics [72], and histopathology [58, 78]. Despite their promise, FL studies often omit key DP parameters, complicating reproducibility.
While FL offers inherent privacy through data decentralization, centralized DP-SGD remains vital for models intended for public sharing or external evaluation. The versatility of DP-SGD across both paradigms underscores its relevance to privacy-preserving clinical AI.
5.1.3 Mechanisms and Privacy Accounting for DP-SGD
@@
The Gaussian mechanism (GM), paired with Rényi differential privacy (RDP), is the dominant configuration in DP-SGD implementations, offering favorable composition properties and practical support through libraries like Opacus and TensorFlow Privacy [85]. Most reviewed works adopt GM with RDP for per-sample gradient perturbation.
Nonetheless, reporting practices are inconsistent. Some studies provide complete details—including $(\epsilon,\delta)$ values, clipping norms, and noise multipliers [13, 5, 66, 75, 81]—while others omit the accounting mechanism or final budget [80, 67, 69]. A few exceptions use the Laplace mechanism (LM) [63] or combine GM and LM [65]. The moments accountant (MA) is used in some centralized and FL settings [83, 73].
Privacy budgets range from conservative ($\epsilon\approx 1$) to more relaxed ($\epsilon\approx 10$). However, few studies systematically explore multiple $\epsilon$ levels, limiting our understanding of utility degradation under stronger privacy guarantees.
Overall, while GM and RDP dominate, standardization in reporting mechanisms, parameters, and accounting tools remains a critical need for enabling rigorous comparisons and reproducibility across studies.
5.2 Design Choices for Privacy-Utility Tradeoffs
@@
Adnan et al. [58] Lung cancer detection MEM-MIL NA NA
Beaulieu-Jones et al. [70] Time series AC-GAN [86] NA NA
Chin-Cheong et al. [61] EHR Wasserstein GAN MSN NA
Chilukoti et al. [60] Histopathology NF-Net [87] NA NA
Daum et al. [62] Cardiac cine MRI (4D) Latent diffusion models[88] NA NA
Fan et al. [66] Tabular data MLP NA NA
Guo, Yu et al. [68] Dermoscopic images NA GN NA
Kaess et al. [71] CT ResNet-9 Scale norm NA
Khanna et al. [72] Genomic data MLP NA NA
Mehmood et al. [74] MRI InceptionV3 [89] NA Affine
Mueller et al. [75] Image-derived graph Graph neural networks Instance norm NA
Nguyen et al.[67] Chest X-ray GAN NA NA
Riess et al. [39] MRI NA GN NA
Tayebi Arasteh et al. [13] Chest X-ray ResNet-9 GN None
Tayebi Arasteh et al. [5] Chest X-ray, abdominal CT ResNet-9 [90] GN None
Tayebi Arasteh et al. [81] Pathological speech ResNet-18 GN None
Yan et al. [64] Chest X-ray ResNet-18 None NA
Zhang et al. [63] Genomic (DNA-SNPs) Custom CNN NA NA
Ziller et al. [84] Chest X-ray, Liver CT VGG-11, Custom U-Net None Affine
@@ 5.2.1 DL Architectures for DP-SGD
Architecture selection significantly influences a model’s resilience to the noise and clipping operations imposed by DP-SGD. Compact CNNs such as ResNet-9 and ResNet-18 were the most widely adopted across classification tasks in chest X-ray, CT, and speech data [13, 5, 64, 81, 71]. Other lightweight backbones like VGG-11 and custom U-Net variants were applied in medical imaging tasks [84], while NF-Net was used in histopathology [60], and InceptionV3 in brain MRI [74].
@@
Generative models used GAN architectures, including Wasserstein GANs for EHR [61], AC-GANs for time series [70], and standard GANs in FL imaging tasks [67]. More recent approaches explored latent diffusion models for cardiac MRI synthesis [62]. MLPs remain prevalent for tabular and genomic data [66, 72], while specialized networks such as graph neural networks (GNNs) [75] and MEM-MIL [58] were applied in niche contexts like graph classification and histopathology.
Notably, vision transformers (ViTs) [91] and other large-scale models were absent from this set, echoing findings in separate comparisons that such architectures are more sensitive to DP noise and tend to underperform under strong privacy constraints.
5.2.2 Normalization Techniques for DP-SGD
Batch normalization (BN), which depends on mini-batch statistics, is generally avoided in DP-SGD because it interferes with per-sample gradient computation. Among studies reporting normalization explicitly, group normalization (GN) was most common, used in both classification and speech tasks [13, 5, 81, 68, 39]. GN avoids batch dependencies and offers more stable training under DP noise.
Other strategies include instance normalization for graph neural networks [75], mode-specific normalization (MSN) for generative EHR models [61], and scale normalization in CT classification [71]. Several studies—particularly those using GANs, MLPs, or CNNs in tabular or genomic settings—did not specify any normalization approach [70, 63, 66, 67, 72, 58]. Others reported disabling normalization entirely [64, 84], likely to avoid incompatibility with gradient clipping.
Overall, there is no universal best practice, though GN appears robust across diverse tasks. Systematic benchmarking of normalization techniques under DP constraints remains a key direction for future research.
5.2.3 Data Augmentation Practices for DP-SGD
Data augmentation, a staple of conventional DL, is rarely applied under DP-SGD. Among reviewed studies, only two used basic affine transformations—one for brain MRI classification [74] and one for multi-modal segmentation [84]. Most others either explicitly avoided augmentation [13, 5, 81] or did not report it.
This reluctance stems from the risk of increasing gradient variance, which exacerbates training instability under noisy DP updates. Augmentations can alter input distributions in ways that interfere with the gradient signal, especially in small medical datasets with high class imbalance [5].
The lack of systematic evaluation limits conclusions about which augmentations, if any, are compatible with DP-SGD. There is a clear need for controlled experiments testing standard and domain-specific augmentation strategies under DP conditions.
5.2.4 Pretraining and Transfer Learning for DP-SGD
Pretraining is a consistent and effective strategy for mitigating performance loss under DP-SGD. By starting from pretrained weights, models avoid the fragile convergence dynamics associated with noisy training from scratch.
@@ 5.2.4 Pretraining and Transfer Learning for DP-SGD
Tayebi Arasteh et al. [5] demonstrated that DP-SGD models trained from scratch failed to learn meaningful features, whereas pretrained ResNet-9 models retained high accuracy. In a related study, pretrained models on chest X-ray and CT were shown to maintain near non-private performance under tight $\epsilon$ budgets [5]. Pretraining on general-domain speech corpora also improved DP training on pathological speech classification [81], despite cross-domain mismatch. Similarly, ImageNet-initialized models improved performance in brain MRI classification under DP [74].
Pretraining is not only helpful but often necessary for training deeper or more expressive models under DP-SGD. Explicit reporting and standardization of pretraining strategies is vital for reproducibility and benchmarking.
5.3 Privacy-Utility Tradeoff Results
A core challenge in differentially private DL is navigating the trade-off between model utility and privacy. As summarized in Table 3, studies vary widely in how they report performance under different privacy budgets, with only a subset evaluating multiple $\epsilon$ levels. This inconsistency limits comparative understanding across clinical modalities and tasks.
@@ 5.3 Privacy-Utility Tradeoff Results
Adnan et al. [58] Histopathology Accuracy = $82\%$ $\epsilon=2.9$/Accuracy = $81\%$ $\epsilon=10$/Accuracy = $78\%$

### v2

@@ Results
Kaess et al. [88FairandPrivateCTContrastAgentDetection] Chest CT Centralized Classification GM/NA
Khanna et al. [32khanna2022privacy] Genomic data FL Classification GM/RDP
Lal et al. [51DeepLearningClassificationofFetal] Tabular data Centralized Classification GM/MA
Mehmood et al. [47mehmood2024balancing] Brain MRI FL Classification GM/NA
Mueller et al. [46DifferentiallyPrivateGraphNeuralNetworksforWhole-GraphClassification] Image-derived graph Centralized Classification GM/RDP
Odeyomi et al. [48PreservingMedicalDatawithRenyiDifferentialPrivacy] Chest X-ray Centralized Classification GM/RDP
P3SGD [64wu2019p3sgd] Histopathology images Centralized Classification GM/MA
Pan et al. [63FedDP] Histopathology images FL Segmentation GM/NA
PP-LDG [59PP-LDG] Abdominal CT Centralized Segmentation GM/MA
Shiri et al. [3DPFRCovid19] Chest CT FL Regression GM/NA
Tayebi Arasteh et al. [13TayebiDomainTransfer] Chest X-ray Centralized Classification GM/RDP
Tayebi Arasteh et al. [22tayebi2024preserving] Chest X-ray, abdominal CT Centralized Classification GM/RDP
Tayebi Arasteh et al. [72arasteh2024differentialprivacyenablesfair] Pathological speech Centralized Classification GM/RDP
Torfi et al. [7RDPCGAN] EEG, EHR, ECG, biopsy Centralized Tabular data generation GM/RDP
Wang et al. [28wang2022privacy] Tabular data FL Classification GM/MA
Ziller et al. [35ziller2021medical] Chest X-ray, abdominal CT, fundoscopy Centralized Classification, segmentation GM/GDP
@@
The reviewed studies apply DP-SGD across a broad range of medical data types, with imaging emerging as the most common modality. Chest X-rays [wang2017chestx] dominate classification tasks [13TayebiDomainTransfer, 22tayebi2024preserving, 24DP-SSLoRA, 30nguyen2021federated, 35ziller2021medical, 50DPFedSAM-Meas, 59PP-LDG, 88FairandPrivateCTContrastAgentDetection], while MRI, particularly brain and cardiac cine MRI, is used in both classification and generative modeling applications [47mehmood2024balancing, 71OnDifferentiallyPrivate3DMedicalImageSynthesiswithControllableLatentDiffusionModels]. Histopathology appears in several studies focusing on cancer classification and segmentation [36adnan2022federated, 63FedDP, 64wu2019p3sgd, 86chilukoti2025differentiallyprivatefinetunednfnet]. Beyond imaging, DP-SGD has been applied to structured data such as EHRs and tabular datasets [26fan2023mitigating, 51DeepLearningClassificationofFetal, 74chin2020generation, 84DifferentiallyPrivateMedicalTextsGenerationUsingGenerativeNeuralNetworks], genomic sequences [25zhang2022secure, 32khanna2022privacy], biosignals including EEG and ECG [7RDPCGAN], and pathological speech [72arasteh2024differentialprivacyenablesfair]. Time series data is also explored in classification contexts [21], and some generative frameworks integrate multimodal data such as EEG, biopsy, and EHR [7RDPCGAN].
@@
Classification remains the predominant downstream task [13TayebiDomainTransfer, 22tayebi2024preserving, 24DP-SSLoRA, 30nguyen2021federated, 35ziller2021medical, 50DPFedSAM-Meas, 88FairandPrivateCTContrastAgentDetection, 47mehmood2024balancing, 86chilukoti2025differentiallyprivatefinetunednfnet, 26fan2023mitigating, 51DeepLearningClassificationofFetal, 32khanna2022privacy, 53Adifferentialprivacybasedprototypicalnetworkformedicaldatalearning, 21], followed by generative modeling [71OnDifferentiallyPrivate3DMedicalImageSynthesiswithControllableLatentDiffusionModels, 74chin2020generation, 84DifferentiallyPrivateMedicalTextsGenerationUsingGenerativeNeuralNetworks, 7RDPCGAN]. Despite growing interest, applications involving longitudinal data or multimodal fusion remain underexplored.
DP-SGD is employed in both centralized [13TayebiDomainTransfer, 22tayebi2024preserving, 24DP-SSLoRA, 25zhang2022secure, 26fan2023mitigating, 35ziller2021medical, 46DifferentiallyPrivateGraphNeuralNetworksforWhole-GraphClassification, 48PreservingMedicalDatawithRenyiDifferentialPrivacy, 51DeepLearningClassificationofFetal, 53Adifferentialprivacybasedprototypicalnetworkformedicaldatalearning, 59PP-LDG, 64wu2019p3sgd, 71OnDifferentiallyPrivate3DMedicalImageSynthesiswithControllableLatentDiffusionModels, 72arasteh2024differentialprivacyenablesfair, 74chin2020generation, 84DifferentiallyPrivateMedicalTextsGenerationUsingGenerativeNeuralNetworks, 86chilukoti2025differentiallyprivatefinetunednfnet, 88FairandPrivateCTContrastAgentDetection, 7RDPCGAN, 21] and federated [3DPFRCovid19, 28wang2022privacy, 30nguyen2021federated, 32khanna2022privacy, 33hatamizadeh2023gradient, 36adnan2022federated, 47mehmood2024balancing, 50DPFedSAM-Meas, 63FedDP] training pipelines. In centralized settings, models are trained on institutional data and privatized using DP-SGD, protecting against post-deployment leakage. These include imaging [13TayebiDomainTransfer, 22tayebi2024preserving, 24DP-SSLoRA, 35ziller2021medical], structured/tabular data [26fan2023mitigating, 51DeepLearningClassificationofFetal], and EHRs [74chin2020generation, 84DifferentiallyPrivateMedicalTextsGenerationUsingGenerativeNeuralNetworks].
@@
FL-based applications involve multi-institutional collaboration, where local clients update models on-site and share only privatized gradients. These setups target imaging [3DPFRCovid19, 30nguyen2021federated, 33hatamizadeh2023gradient], genomics [32khanna2022privacy], and histopathology [36adnan2022federated, 63FedDP]. Despite their promise, FL studies often omit key DP parameters, complicating reproducibility. While FL offers inherent privacy through data decentralization [tayebi2023collaborative, tayebi2023enhancing, FederatedLearningSource], centralized DP-SGD remains vital for models intended for public sharing or external evaluation. The versatility of DP-SGD across both paradigms underscores its relevance to privacy-preserving clinical AI.
@@
The Gaussian mechanism (GM), paired with RDP, is the dominant configuration in DP-SGD implementations, offering favorable composition properties and practical support through libraries like Opacus [opacus_neurips2021] and TensorFlow Privacy [dahl2018private, papernot2019machine]. Most reviewed works adopt GM with RDP for per-sample gradient perturbation.
Nonetheless, reporting practices are inconsistent. Some studies provide complete details, including $(\epsilon,\delta)$ values, clipping norms, and noise multipliers [13TayebiDomainTransfer, 22tayebi2024preserving, 26fan2023mitigating, 46DifferentiallyPrivateGraphNeuralNetworksforWhole-GraphClassification, 72arasteh2024differentialprivacyenablesfair], while others omit the accounting mechanism or final budget [3DPFRCovid19, 30nguyen2021federated, 33hatamizadeh2023gradient]. A few exceptions use the Laplace mechanism (LM) [25zhang2022secure] or combine GM and LM [50DPFedSAM-Meas]. The moments accountant (MA) is used in some centralized and FL settings [28wang2022privacy, 51DeepLearningClassificationofFetal]. Privacy budgets range from conservative ($\epsilon\approx 1$) to more relaxed ($\epsilon\approx 10$). However, few studies systematically explore multiple $\epsilon$ levels, limiting our understanding of utility degradation under stronger privacy guarantees. Overall, while GM and RDP dominate, standardization in reporting mechanisms, parameters, and accounting tools remains a critical need for enabling rigorous comparisons and reproducibility across studies.
Figure 4: Integrating differential privacy (DP) into the lifecycle of medical deep learning models. (a) Institutions curate data and remove direct identifiers or metadata before training begins. (b) DP can be applied in multiple configurations, including centralized DP, where all data are stored on a single server and DP-SGD is applied during training, federated learning (FL) with server-side DP, where institutions keep data locally, send model updates to a central server, and noise is added at the server before aggregation, and FL with local DP, where noise is added at each institution before sharing updates, providing stronger protection against an untrusted server. (c) The trained model is distributed to hospitals or clinical sites for inference. (d) After deployment, model behavior is monitored for privacy and security risks, including anomaly detection, privacy-leakage checks (e.g., membership inference signals), and tracking cumulative privacy budget when repeated queries or fine-tuning occur. Representative chest X-ray images are provided by the ChestX-ray14 dataset from NIH Clinical Center [wang2017chestx].
Design choices for privacy-utility tradeoffs
@@
Adnan et al. [36adnan2022federated] Histopathology images MEM-MIL NA NA
Beaulieu-Jones et al. [21] Time series AC-GAN [odena2017conditional] NA NA
Chin-Cheong et al. [74chin2020generation] EHR Wasserstein GAN MSN NA
Chilukoti et al. [86chilukoti2025differentiallyprivatefinetunednfnet] Histopathology NF-Net [brock2021high] NA NA
Daum et al. [71OnDifferentiallyPrivate3DMedicalImageSynthesiswithControllableLatentDiffusionModels] Cardiac cine MRI Latent diffusion models [rombach2022high] NA NA
Fan et al. [26fan2023mitigating] Tabular data MLP NA NA
Guo, Yu et al. [53Adifferentialprivacybasedprototypicalnetworkformedicaldatalearning] Dermoscopic photography NA GN NA
Kaess et al. [88FairandPrivateCTContrastAgentDetection] Chest CT ResNet-9 [he2016deep] Scale norm NA
Khanna et al. [32khanna2022privacy] Genomic data MLP NA NA
Mehmood et al. [47mehmood2024balancing] Brain MRI InceptionV3 [szegedy2015going] NA Affine
Mueller et al. [46DifferentiallyPrivateGraphNeuralNetworksforWhole-GraphClassification] Image-derived graph Graph neural networks Instance norm NA
Nguyen et al.[30nguyen2021federated] Chest X-ray GAN NA NA
Riess et al. [73Complex-ValuedFederatedLearningwithDifferentialPrivacy] Brain MRI NA GN NA
Tayebi Arasteh et al. [13TayebiDomainTransfer] Chest X-ray ResNet-9 GN None
Tayebi Arasteh et al. [22tayebi2024preserving] Chest X-ray, abdominal CT ResNet-9 GN None
Tayebi Arasteh et al. [72arasteh2024differentialprivacyenablesfair] Pathological speech ResNet-18 GN None
Yan et al. [24DP-SSLoRA] Chest X-ray ResNet-18 None NA
Zhang et al. [25zhang2022secure] Genomic data Custom CNN NA NA
Ziller et al. [35ziller2021medical] Chest X-ray, abdominal CT VGG-11, custom U-Net None Affine
@@
Architecture selection significantly influences a model’s resilience to the noise and clipping operations imposed by DP-SGD. Compact CNNs such as ResNet-9 and ResNet-18 [he2016deep] were the most widely adopted across classification tasks in chest X-ray, CT, and speech data [13TayebiDomainTransfer, 22tayebi2024preserving, 24DP-SSLoRA, 72arasteh2024differentialprivacyenablesfair, 88FairandPrivateCTContrastAgentDetection]. Other lightweight backbones like VGG-11 and custom U-Net variants were applied in medical imaging tasks [35ziller2021medical], while NF-Net was used in histopathology [86chilukoti2025differentiallyprivatefinetunednfnet], and InceptionV3 in brain MRI [47mehmood2024balancing].
@@
Generative models used GAN architectures, including Wasserstein GANs for EHR [74chin2020generation], AC-GANs for time series [21], and standard GANs in FL imaging tasks [30nguyen2021federated]. More recent approaches explored latent diffusion models for cardiac MRI synthesis [71OnDifferentiallyPrivate3DMedicalImageSynthesiswithControllableLatentDiffusionModels]. MLPs remain prevalent for tabular and genomic data [26fan2023mitigating, 32khanna2022privacy], while specialized networks such as graph neural networks (GNNs) [46DifferentiallyPrivateGraphNeuralNetworksforWhole-GraphClassification] and MEM-MIL [36adnan2022federated] were applied in niche contexts like graph classification and histopathology. Notably, vision transformers (ViTs) [mainvit] and other large-scale models were absent from this set, echoing findings in separate comparisons that such architectures are more sensitive to DP noise and tend to underperform under strong privacy constraints.
Batch normalization (BN) [ioffe2015batch], which depends on mini-batch statistics, is generally avoided in DP-SGD because it interferes with per-sample gradient computation [22tayebi2024preserving, kaissis2020secure, 13TayebiDomainTransfer, DPSGD, 72arasteh2024differentialprivacyenablesfair]. Among studies reporting normalization explicitly, group normalization (GN) [wu2018group] was most common, used in both classification and speech tasks [13TayebiDomainTransfer, 22tayebi2024preserving, 72arasteh2024differentialprivacyenablesfair, 53Adifferentialprivacybasedprototypicalnetworkformedicaldatalearning, 73Complex-ValuedFederatedLearningwithDifferentialPrivacy]. GN avoids batch dependencies and offers more stable training under DP noise. Other strategies include instance normalization for graph neural networks [46DifferentiallyPrivateGraphNeuralNetworksforWhole-GraphClassification], mode-specific normalization (MSN) [deecke2018mode] for generative EHR models [74chin2020generation], and scale normalization in CT classification [88FairandPrivateCTContrastAgentDetection]. Several studies, particularly those using GANs, MLPs, or CNNs in tabular or genomic settings, did not specify any normalization approach [21, 25zhang2022secure, 26fan2023mitigating, 30nguyen2021federated, 32khanna2022privacy, 36adnan2022federated]. Others reported disabling normalization entirely [24DP-SSLoRA, 35ziller2021medical], likely to avoid incompatibility with gradient clipping. Overall, there is no universal best practice, though GN appears robust across diverse tasks. Systematic benchmarking of normalization techniques under DP constraints remains a key direction for future research.
Data augmentation, a staple of conventional DL, is rarely applied under DP-SGD. Among reviewed studies, only two used basic affine transformations, one for brain MRI classification [47mehmood2024balancing] and one for multi-modal segmentation [35ziller2021medical]. Most others either explicitly avoided augmentation [13TayebiDomainTransfer, 22tayebi2024preserving, 72arasteh2024differentialprivacyenablesfair] or did not report it. This reluctance stems from the risk of increasing gradient variance, which exacerbates training instability under noisy DP updates. Augmentations can alter input distributions in ways that interfere with the gradient signal, especially in small medical datasets with high class imbalance [22tayebi2024preserving]. The lack of systematic evaluation limits conclusions about which augmentations, if any, are compatible with DP-SGD. There is a clear need for controlled experiments testing standard and domain-specific augmentation strategies under DP conditions.
@@
Pretraining is a consistent and effective strategy for mitigating performance loss under DP-SGD [13TayebiDomainTransfer, 22tayebi2024preserving, 72arasteh2024differentialprivacyenablesfair]. By starting from pretrained weights, models avoid the fragile convergence dynamics associated with noisy training from scratch. Tayebi Arasteh et al. [22tayebi2024preserving] demonstrated that DP-SGD models trained from scratch failed to learn meaningful features, whereas pretrained ResNet-9 models retained high accuracy. In a related study, pretrained models on chest X-ray and CT were shown to maintain near non-private performance under tight $\epsilon$ budgets [22tayebi2024preserving]. Pretraining on general-domain speech corpora also improved DP training on pathological speech classification [72arasteh2024differentialprivacyenablesfair], despite cross-domain mismatch. Similarly, ImageNet-initialized models improved performance in brain MRI classification under DP [47mehmood2024balancing]. Pretraining is not only helpful but often necessary for training deeper or more expressive models under DP-SGD [72arasteh2024differentialprivacyenablesfair]. Explicit reporting and standardization of pretraining strategies are vital for reproducibility and benchmarking.
A core challenge in differentially private DL is navigating the tradeoff between model utility and privacy [22tayebi2024preserving, kaissis2020secure]. As summarized in Table 3, studies vary widely in how they report performance under different privacy budgets, with only a subset evaluating multiple $\epsilon$ levels. This inconsistency limits comparative understanding across clinical modalities and tasks. At very low privacy budgets ($\epsilon\approx 1$), the magnitude of injected noise often leads to significant performance degradation. For example, Tayebi Arasteh et al. [22tayebi2024preserving] observed an AUC drop from 89.7% to 84.0% on chest X-rays and from 99.7% to 92.0% on abdominal CT when applying $\epsilon=0.5$. In pathological speech classification, accuracy fell from 99.1% to 88.3% at $\epsilon=0.9$ [72arasteh2024differentialprivacyenablesfair], and in dermoscopy, accuracy declined from 69.1% to  43% at $\epsilon=3$ [53Adifferentialprivacybasedprototypicalnetworkformedicaldatalearning]. Similar trends were reported for generative tasks, such as cardiac MRI synthesis, where Fréchet inception distance (FID) worsened substantially under $\epsilon=1$ [71OnDifferentiallyPrivate3DMedicalImageSynthesiswithControllableLatentDiffusionModels]. However, there are exceptions. In a multi-institutional study on domain transfer, Tayebi Arasteh et al. [13TayebiDomainTransfer] showed that external validation performance remained nearly unchanged across $\epsilon\approx 1$–10. This suggests that DP noise may act as a regularizer, benefiting generalization even as in-domain performance drops.
Most reviewed studies operate in the strong privacy regime ($\epsilon\approx 5$–10), where performance generally remains close to non-private baselines. For instance, chest X-ray AUC dropped modestly from 89.7% to 87.4% at $\epsilon=7.9$, and CT classification remained above 99% even at $\epsilon=8$ [22tayebi2024preserving]. In MRI [47mehmood2024balancing] and genomic data [32khanna2022privacy], performance losses at $\epsilon=10$ were similarly modest. Nonetheless, more fragile domains show greater vulnerability to DP noise. Studies in histopathology [36adnan2022federated], dermoscopy [53Adifferentialprivacybasedprototypicalnetworkformedicaldatalearning], and speech [72arasteh2024differentialprivacyenablesfair] report larger performance drops, reflecting higher input variance, smaller datasets, and more complex learning tasks. Tabular and sensor-based applications, e.g., in mental health or activity monitoring, also showed substantial utility loss at lower $\epsilon$ values [31AZIZ2022104113, 60DifferentialPrivateFederatedTransferLearningforMentalHealthMonitoringinEverydaySettings]. These trends emphasize the importance of modality-specific tuning [72arasteh2024differentialprivacyenablesfair]: modalities like chest imaging are more robust under DP-SGD [13TayebiDomainTransfer, 22tayebi2024preserving], while others require additional regularization, pretraining, or architectural adaptation to maintain clinical viability.
While most studies report $\epsilon$, the value of $\delta$, which bounds the probability of privacy failure, is often missing or inconsistently stated. Best practice recommends $\delta\leq 1/n$, where $n$ is the dataset size. Many studies follow this, setting $\delta$ between $10^{-5}$ and $10^{-6}$, but others omit it or use heuristic values without justification.
Reporting gaps extend beyond privacy parameters. Performance under multiple $\epsilon$ levels is rarely benchmarked, and core implementation details, such as clipping norm, noise multiplier, accounting method (e.g., RDP or moments accountant [DPSGD]), and number of training epochs, are frequently missing. This limits reproducibility and cross-study comparison. This issue is particularly acute in FL settings, where DP is often applied to distributed model updates. Although FL introduces additional privacy through decentralization [tayebi2023collaborative, tayebi2023enhancing, FederatedLearningSource, tayebiarasteh23_interspeech], few studies evaluate how DP noise affects utility across varying budgets in such settings, hindering generalizability.
To address these gaps, we recommend several reporting standards for future work. (i) Researchers should report both non-private and private model performance to provide clear context for evaluating the impact of privacy mechanisms. (ii) The values of key privacy parameters, including $\epsilon$, $\delta$, and the accounting method used (e.g., RDP or moments accountant), should be explicitly stated to ensure interpretability and comparability. (iii) Performance metrics should be reported across multiple $\epsilon$ levels to characterize the tradeoff between privacy and utility. (iv) Finally, studies should document core model and training details, including architecture type, dataset size, pretraining strategies, and other relevant hyperparameters. Adopting these practices will improve transparency, foster reproducibility, and support more rigorous benchmarking of privacy-preserving models for clinical deployment.
@@
Adnan et al. [36adnan2022federated] Histopathology images Accuracy = $82\%$ $\epsilon=2.9$/Accuracy = $81\%$ $\epsilon=10$/Accuracy = $78\%$
Al Aziz et al. [31AZIZ2022104113] Tabular data AUC = $98\%$ $\epsilon=1/$AUC = $70\%$ $\epsilon=10/$AUC = $92\%/$
Alsenani et al. [23FAItH] Sensor data Accuracy = $30\%$ NA $\epsilon=10$/Accuracy = $27\%$
Chin-Cheong et al. [74chin2020generation] EHR AUC = $80.0\%$ $\epsilon=1$/AUC = $66.1\%$ $\epsilon=10$/AUC = $67.8\%$
Daum et al. [71OnDifferentiallyPrivate3DMedicalImageSynthesiswithControllableLatentDiffusionModels] Cardiac cine MRI FID = $15.4$ $\epsilon=1$/FID = $29.8$ $\epsilon=10$/FID = $26.8$
Fan et al. [26fan2023mitigating] Tabular data IBS = ~$17$ $\epsilon=1$/IBS = ~$21$ $\epsilon=8$/IBS = ~$18$
Fu et al. [82DP-MLD] EEG, sensor data Accuracy = $99.3\%$ $\epsilon=1$/Accuracy = $98.7\%$ NA
Gong et al. [27gong2023federated] Time series Accuracy = $88.5\%$ $\epsilon=2$/Accuracy = $75.3\%$ $\epsilon=10$/Accuracy = $78.2\%$
Guo, Yu et al. [53Adifferentialprivacybasedprototypicalnetworkformedicaldatalearning] Dermoscopic photography Accuracy = $69.1\%$ $\epsilon=3$/Accuracy = $43\%$ $\epsilon=10$/Accuracy = $61.2\%$
Kaess et al. [88FairandPrivateCTContrastAgentDetection] Chest CT Accuracy = $99.5\%$ NA $\epsilon=8$ /Accuracy = $97.4\%$
Mehmood et al. [47mehmood2024balancing] Brain MRI Accuracy = $88\%$ NA $\epsilon=5.6$/Accuracy = $78\%$
Ming Y et al. [42GWON2024107738] Tabular EMR DWS =$0.0041$ $\epsilon=1$/DWS = $0.05$ $\epsilon=10$/DWS = $0.0061$
Mueller et al. [46DifferentiallyPrivateGraphNeuralNetworksforWhole-GraphClassification] Image-derived graph AUC = $98.5\%$ $\epsilon=1$/AUC = $95.1\%$ $\epsilon=10$/AUC = $97.2\%$
Nguyen et al. [30nguyen2021federated] Chest X-ray Accuracy = $99\%$ $\epsilon=0.3$/Accuracy = $90\%$ NA
Odeyomi et al. [49yuan2019collaborative] Chest X-ray Accuracy = $92\%$ $\epsilon=2$/Accuracy = $88\%$ $\epsilon=8$/Accuracy = $90\%$
Riess et al. [73Complex-ValuedFederatedLearningwithDifferentialPrivacy] Brain MRI Accuracy = $90.1\%$ $\epsilon=1$/Accuracy = $81.6\%$ $\epsilon=10$/Accuracy = $90\%$
Sun et al. [39SUN2023104404] Tabular data DOP = ~$0.75$ $\epsilon=1$/DOP = ~$1.2$ $\epsilon=10$/DOP = ~$1.2$
Tang et al. [34tang2024personalized] Tabular data F1-score = $19.6\%$ $\epsilon=1$/F1-score = $19.79\%$ NA